# WaveFormer: A Lightweight Transformer Model for sEMG-based Gesture Recognition

### v1

@@ Abstract
Human-machine interaction, particularly in prosthetic and robotic control, has seen progress with gesture recognition via surface electromyographic (sEMG) signals. However, classifying similar gestures that produce nearly identical muscle signals remains a challenge, often reducing classification accuracy. Traditional deep learning models for sEMG gesture recognition are large and computationally expensive, limiting their deployment on resource-constrained embedded systems. In this work, we propose WaveFormer, a lightweight transformer-based architecture tailored for sEMG gesture recognition. Our model integrates time-domain and frequency-domain features through a novel learnable wavelet transform, enhancing feature extraction. In particular, the WaveletConv module, a multi-level wavelet decomposition layer with depthwise separable convolution, ensures both efficiency and compactness. With just 3.1 million parameters, WaveFormer achieves 95% classification accuracy on the EPN612 dataset, outperforming larger models. Furthermore, when profiled on a laptop equipped with an Intel CPU, INT8 quantization achieves real‑time deployment with a 6.75 ms inference latency.
Clinical relevance— Accurate decoding of fine finger movements from forearm sEMG enables more natural multi‑digit control in myoelectric prostheses and gives therapists an objective tool for tracking hand‑function recovery.
@@ I INTRODUCTION
Surface electromyographic (sEMG) signals are a rich, non‑invasive proxy for muscle activity and support applications ranging from prosthetic control to rehabilitation and gesture‑based human-computer interaction. Classical sEMG pipelines rely on hand-crafted descriptors that include characteristics of the time domain such as mean absolute value, zero crossings, and waveform length, together with basic spectral indices, which are then fed to shallow classifiers such as $k$ NN, LDA or SVMs [1, 2]. Although simple, this scheme has three well‑known limitations. First, performance is fragile: Even minor electrode shifts or skin impedance changes can dramatically degrade precision because descriptors are highly sensitive to spatial sampling and session‑to‑session drift [3, 4]. Second, representing each window with only a few dozen engineered numbers constrains the expressive power of the model, particularly when faced with larger and more diverse datasets that exhibit bursty, non‑stationary patterns and inter‑channel correlations [5, 6]. Third, competitive results require laborious adjustment of window lengths, feature subsets, and classifier hyper‑parameters, followed by per‑subject recalibration, which impedes broader adoption [7, 8]. These shortcomings have spurred interest in end‑to‑end deep architectures that learn robust, multi-resolution representations directly from raw sEMG recordings. Recent advances in deep learning have shown that these models can outperform classical methods, offering new possibilities for real-time applications [9, 10].
@@ II-A Surface Electromyographic Signals
[…]
@@ II-B Wavelet Transform and RoPEAttention
Figure 1: Overview of WaveFormer. The raw sEMG input $\mathbf{X}\in\mathbb{R}^{C\times T}$ is first segmented into time-domain patches via a learnable patch embedding. These patches are then fed into a WaveletConv module that applies multi-level wavelet decompositions and reconstructions to extract rich, multi-scale features. Next, the resulting representations are passed into a Transformer encoder equipped with RoPEAttention to capture global temporal correlations. Finally, a classification head produces likelihood scores for each hand gesture class, which are used to predict the corresponding gesture.
@@
In practical sEMG applications, the DWT’s multilevel filter–bank implementation splits the raw signal into four two-dimensional subbands at every scale. The low-low (LL) branch, obtained by low-pass filtering across both channel and time axes, stores the slowly varying envelope that reflects general muscle activation trends, and this approximation is recursively decomposed to reveal still coarser behavior. The low–high (LH) band retains rapid temporal transients that are common to all channels, capturing brief motor‑unit bursts; the high–low (HL) band emphasizes spatial differences that change slowly over time, thus disentangling interelectrode crosstalk or electrode placement effects; and the high–high (HH) band isolates short-lived channel-specific high-frequency spikes, including high-frequency noise. Taken together, this coarse‑to‑fine hierarchy preserves global activation patterns while simultaneously exposing localized temporal and spatial details, providing a noise-resistant representation that has proven effective for gesture recognition, onset detection, and artifact suppression under diverse contraction intensities and recording conditions [16, 17].
@@ III METHODS
III-A Patch Embedding and 2D Feature Representation
Consider an sEMG segment $\mathbf{X}\in\mathbb{R}^{C\times T}$ with $C$ channels and $T$ time samples. The input is cropped or zero-padded to a fixed length $T$. As illustrated in Fig. 1, the signal is divided into non-overlapping patches using a 2D convolution operation. Specifically, the patch embedding module applies a 2D convolution with kernel size $(1,P)$ and stride $(1,P)$, where $P$ is the patch width along the temporal dimension. This operation simultaneously partitions the signal into $N=T/P$ patches per channel and maps each patch from its original $P$ dimensional representation to an embedding of dimension $D$. The convolution is followed by LayerNorm and GELU activation, yielding a 2D feature map of shape $(D,C,N)$ where $D$ is the embedding dimension, $C$ is the number of channels, and $N$ is the number of temporal patches. This representation preserves both spatial (channel) and temporal relationships. This 2D representation enables subsequent wavelet convolution to perform multiscale decomposition along both channel and time dimensions, while maintaining a fine-grained temporal context for efficient transformer processing.
III-B WaveletConv
@@
III-C Transformer-Based Classification via RoPEAttention
@@
where $y_{c}$ represents one-hot labels and $\sigma(\cdot)$ is the softmax function. During inference, the predicted gesture corresponds to $\arg\max(\hat{\mathbf{y}})$.
@@ IV EXPERIMENTAL RESULTS
TABLE I: Performance comparison (Acc., F1, AUROC) on four sEMG datasets. The best value in each column is shown in bold.
@@ IV-A Data and Preprocessing
To thoroughly evaluate our method, we utilize four publicly available sEMG datasets that offer various conditions and sensor configurations. The EPN‑612 dataset comprises 612 subjects who perform five hand gestures (wave‑in, wave‑out, pinch, open, fist) plus a relaxed state, with each subject providing 50 samples per gesture [24]. The Ninapro DB5 dataset includes ten intact subjects who repeat 52 hand movements plus rest, acquired at 200 Hz using two Thalmic Myo armbands [25]. The Ninapro DB6 focuses on repeatability between sessions, with ten subjects recorded in 2 kHz using 14 Delsys Trigno electrodes, performing seven grasps over five days [26]. The UC Irvine EMG dataset provides raw 200 Hz sEMG of 36 subjects who perform static hand gestures for 3 s intervals [27].
We apply a unified preprocessing pipeline across all datasets. Channel-wise Z-score normalization is performed to reduce intersubject variability and sensor artifacts. Subsequently, we segment the continuous signals into fixed‑length windows of 200 samples with 50% overlap, providing sufficient temporal context while maintaining computational efficiency. For patch embedding, we set the patch width to 40, creating five temporal patches per window that capture local muscle activation patterns.
@@ IV-E Deployment Efficiency and Model Quantization
INT8 quantization achieves significant speedup with minimal accuracy loss. The quantized model delivers an average latency 6.75 ms per sample - a 34% reduction from the FP32 baseline (10.23 ms). The throughput increases from 97.8 to 148.1 queries per second, representing an improvement 51%. Memory efficiency also improves substantially, with CPU memory usage dropping from 84.9 MB to 65.2 MB. These results demonstrate that WaveFormer, with only 3.10 M parameters, achieves practical deployment efficiency on standard CPU hardware for real‑time sEMG processing systems.

### v2

@@ Abstract
Human-machine interaction, particularly in prosthetic and robotic control, has seen progress with gesture recognition via surface electromyographic (sEMG) signals. However, classifying similar gestures that produce nearly identical muscle signals remains a challenge, often reducing classification accuracy. Traditional deep learning models for sEMG gesture recognition are large and computationally expensive, limiting their deployment on resource-constrained embedded systems. In this work, we propose WaveFormer, a lightweight transformer-based architecture tailored for sEMG gesture recognition. Our model integrates time-domain and frequency-domain features through a novel learnable wavelet transform, enhancing feature extraction. In particular, the WaveletConv module, a multi-level wavelet decomposition layer with depthwise separable convolution, ensures both efficiency and compactness. With just 3.1 million parameters, WaveFormer achieves 95% classification accuracy on the EPN612 dataset, outperforming larger models. Furthermore, when profiled on a laptop equipped with an Intel CPU, INT8 quantization achieves real-time deployment with a 6.75 ms inference latency. Code and data are available at https://github.com/ForeverBlue816/WaveFormer.
@@ I INTRODUCTION
Surface electromyographic (sEMG) signals are a rich, non‑invasive proxy for muscle activity and support applications ranging from prosthetic control to rehabilitation and gesture‑based human-computer interaction. Classical sEMG pipelines rely on hand-crafted descriptors that include characteristics of the time domain such as mean absolute value, zero crossings, and waveform length, together with basic spectral indices, which are then fed to shallow classifiers such as $k$ NN, LDA or SVMs [1]. Although simple, this scheme has three well‑known limitations. First, performance is fragile: Even minor electrode shifts or skin impedance changes can dramatically degrade precision because descriptors are highly sensitive to spatial sampling and session‑to‑session drift [2]. Second, representing each window with only a few dozen engineered numbers constrains the expressive power of the model, particularly when faced with larger and more diverse datasets that exhibit bursty, non‑stationary patterns and inter‑channel correlations [3]. Third, competitive results require laborious adjustment of window lengths, feature subsets, and classifier hyper‑parameters, followed by per‑subject recalibration, which impedes broader adoption [4, 5]. These shortcomings have spurred interest in end‑to‑end deep architectures that learn robust, multi-resolution representations directly from raw sEMG recordings. Recent advances in deep learning have shown that these models can outperform classical methods, offering new possibilities for real-time applications [6, 7].
@@ II-A Surface Electromyographic Signals
Surface EMG (sEMG) records the bioelectric potentials generated by muscle fibers during contraction. Typical amplitudes range from $10\,\mu\mathrm{V}$ to $1\,\mathrm{mV}$, and most control applications require a bandwidth of up to about $2\,\text{kHz}$; specialized motor unit studies can sample at $\sim\!10\,\text{kHz}$[5]. Signals are captured non-invasively using gel- or adhesive-based electrodes placed on the skin above the target muscle. However, measurement quality is highly sensitive to the skin–electrode interface: contact impedance, electrode placement, and adaptation to daily use introduce drift and variability[1]. Further degradation arises from motion artefacts, cable microphonics, and common‑mode interference when the reference ground is unstable. Robust sEMG pipelines therefore devote considerable effort to mitigating these noise sources and normalizing recordings across sessions.
@@ II-B Wavelet Transform and RoPEAttention
Figure 1: Overview of WaveFormer. The raw sEMG input $\mathbf{X}\in\mathbb{R}^{C\times T}$ is first segmented into time-domain patches via a learnable patch projector. These patches are then fed into a WaveletConv module that applies multi-level wavelet decompositions and reconstructions to extract rich, multi-scale features. Next, the resulting representations are passed into a Transformer encoder equipped with RoPEAttention to capture global temporal correlations. Finally, a classification head produces likelihood scores for each hand gesture class, which are used to predict the corresponding gesture.
@@
In practical sEMG applications, the DWT’s multilevel filter–bank implementation splits the raw signal into four two-dimensional subbands at every scale. The low-low (LL) branch, obtained by low-pass filtering across both channel and time axes, stores the slowly varying envelope that reflects general muscle activation trends, and this approximation is recursively decomposed to reveal still coarser behavior. The low–high (LH) band retains rapid temporal transients that are common to all channels, capturing brief motor‑unit bursts; the high–low (HL) band emphasizes spatial differences that change slowly over time, thus disentangling interelectrode crosstalk or electrode placement effects; and the high–high (HH) band isolates short-lived channel-specific high-frequency spikes, including high-frequency noise. Taken together, this coarse‑to‑fine hierarchy preserves global activation patterns while simultaneously exposing localized temporal and spatial details, providing a noise-resistant representation that has proven effective for gesture recognition, onset detection, and artifact suppression under diverse contraction intensities and recording conditions [12].
@@ III METHODS
III-A Preprocessing
Raw sEMG recordings are first filtered and normalized. We apply a zero-phase bandpass filter to suppress motion artifacts and high-frequency noise (e.g., 20-450 Hz, 4th order), followed by a powerline notch filter (50/60 Hz, quality factor Q=30Q{=}30) to remove mains interference. Each channel is then standardized via per-channel $z$-score normalization. The continuous stream is segmented into fixed-length windows using a sliding window (e.g., window length T=1024T{=}1024 samples with 50% overlap), and each window is cropped or zero-padded to length $T$ if necessary. After these steps, each sample is formatted as $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ is the number of electrodes and $T$ the window length.
III-B Patch Embedding and 2D Feature Representation
Consider a preprocessed sEMG segment $\mathbf{X}\in\mathbb{R}^{C\times T}$ with $C$ channels and $T$ time samples. As illustrated in Fig. 1, the signal is first processed by a patch projector, a learnable 2D convolution-based embedding layer, that partitions the signal into non-overlapping patches and maps each to a fixed-dimensional latent vector. Specifically, the patch projector applies a 2D convolution with kernel size $(1,P)$ and stride $(1,P)$, where $P$ is the patch width along the temporal dimension. In our implementation, we use $T=1024$ time samples with patch size $P=64$, resulting in $N=16$ temporal patches. This operation simultaneously divides the signal into N=T/PN{=}T/P patches per channel and projects each $P$-dimensional raw patch into an embedding of dimension $D=256$. The convolution is followed by LayerNorm and GELU activation, yielding a 2D feature map of shape $(D,C,N)$. This representation preserves both spatial (channel) and temporal relationships, enabling the subsequent wavelet convolution to perform multiscale decomposition along both channel and time dimensions, while maintaining a fine-grained temporal context for efficient transformer processing.
III-C WaveletConv
@@
III-D Transformer-Based Classification via RoPEAttention
@@
[…]
@@ IV EXPERIMENTAL RESULTS
[…]
@@ IV-A Data and Preprocessing
To thoroughly evaluate our method, we utilize four publicly available sEMG datasets that offer various conditions and sensor configurations.
EPN‑612 dataset comprises 612 subjects who perform five hand gestures (wave‑in, wave‑out, pinch, open, fist) plus a relaxed state, with each subject providing 50 samples per gesture [19]. Ninapro DB5 dataset includes ten intact subjects who repeat 52 hand movements plus rest, acquired at 200 Hz using two Thalmic Myo armbands [20]. Ninapro DB6 focuses on repeatability between sessions, with ten subjects recorded in 2 kHz using 14 Delsys Trigno electrodes, performing seven grasps over five days [21]. UC Irvine EMG dataset provides raw 200 Hz sEMG of 36 subjects who perform static hand gestures for 3 s intervals [22].
All datasets are pre-processed using standard EMG procedures as explained in III-A, including bandpass filtering, notch filtering, $z$ score normalization, and fixed-length window segmentation, resulting in samples of shape $\mathbf{X}\in\mathbb{R}^{C\times T}$.
@@ IV-E Deployment Efficiency and Model Quantization
INT8 quantization achieves significant speedup with minimal accuracy loss. The quantized model delivers an average latency 6.75 ms per sample, which is a 34% reduction from the FP32 baseline (10.23 ms). The throughput increases from 97.8 to 148.1 queries per second, representing an improvement 51%. Memory efficiency also improves substantially, with CPU memory usage dropping from 84.9 MB to 65.2 MB. These results demonstrate that WaveFormer, with only 3.10 M parameters, achieves practical deployment efficiency on standard CPU hardware for real‑time sEMG processing systems.
@@ V CONCLUSIONS
ETHICAL APPROVAL
The sEMG datasets used in this study (EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG) were collected following approved ethical protocols. The original data collection procedures involving human subjects were approved by the respective Institutional Review Boards of the institutions where the data were collected. All participants provided informed consent, and the experimental procedures followed the principles outlined in the Declaration of Helsinki. No additional human subject experiments were conducted specifically for this study.